OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 8 and 15 units in length.

OpenStudy (anonymous):

\[\sqrt{8^{2}+15^{2}}\]

OpenStudy (anonymous):

\[\sqrt{8^2+15^2}\]

OpenStudy (anonymous):

i get \[\sqrt{289}=17\]

OpenStudy (anonymous):

17

OpenStudy (anonymous):

what hakkanumatata said